The demand for product X depends on the price of product X as well as the average household income (Y) according to the following relationship Qdx = 800 - 35 P + 0.001Y The supply of product X is positively related to own price of product X and negatively dependent upon W, the price of some input. This relationship is expressed as: Qsx = 130 + 30 P - 2 W Given that Y = 50,000 and W = 10, what is the: 1. Equilibrium price? 2. Equilibrium quantity? Suppose that income increases to 60,000 and W remains constant. What is the new: 3. Equilibrium price? 4. Equilibrium quantity? Assuming that income remains constant at 60,000 and W increases to 15, what is the new: 5: Equilibrium price? 6. Equilibrium quantity?
Answer #1

Qdx=800-35P+0.001Y

Qsx=130+30P-2W

1. Y=50000, W=10

Qdx= 800-35P+0.001*50000= 850-35P

Qsx=130+30P-2×10= 110+30P

Equilibrium:

850-35P=110+30P

Equilibrium Price P*= 740/65= $11.38

2. Equilibrium quantity Q*=110+30*11.38=451.4

3. Y=$60000, W=10

Qdx=800-35P+0.001*60000= 860-35P

Qsx= 110+30P

Equilibrium:

860-35P=110+30P

Equilibrium Price P*= 750/65=$11.53

4. Equilibrium quantity Q*= 110+30*11.53=455.9

5. Y=60000, W=15

Qdx=800-35P+0.001*60000= 860-35P

Qsx= 130+30P-2*15= 100+30P

Equilibrium:

860-35P=100+30P

Equilibrium Price P*= 760/65=$11.69

6. Equilibrium quantity Q*= 100+30*11.69= 450.7